Mila is saving for university. She has $9000, of which some is in a savings account earning 3.25% per year and the rest is in a GIC earning 5% a year. If she earned $324 in simple interest last year in total, how much was invested in her GIC?
Answer by mananth(16949)

let her invest $ x in GIC at 5%
$9000-x in savings at 3.25%
interest earned = $324
...
0.05x+0.0325(9000-x)=324
0.05x-0.0325x+292.5 =324
0.0175x=31.5
x= 31.5/0.0175
$1800 in GIC
....
m.ananth@hotmail.ca